# On a combinatorial classification of fine compactified universal Jacobians

### v1

@@ 3.4. Comparison with classical fine compactified universal Jacobians
In this section, we use our classification in terms of universal stability conditions to construct explicit examples of non-classical fine compactified universal Jacobians. Not only, we show that in general, for any $g\geq 1$ the class of fine compactified universal Jacobians is strictly larger than the class of classical fine compactified universal Jacobian described in [KP19], but we also find the minimum number of markings $n$ for this to be true over ¯$\overline{\mathcal{M}}_{g,n}$.
Moreover, we are able to make a further distinction between non-classical fine compactified universal Jacobians whose fibres over geometric points are all classical compactified Jacobians and those for which there exist non-classical fibres. In particular, for genus $g=2$, with the help of computations by Rhys Wells [Wel23], we see that there is a range $4\leq n\leq 5$ where not all the fine compactified universal Jacobians are classical, but their restrictions over all geometric points are.
In the body of [PT23, Section $10$] Question $\star$1 is asked. A complete answer is given in Part $1$ of Theorem B, which will be proven via the examples and propositions of this section, together with the following
@@
$\begin{Bmatrix}\textup{Fine classical compactified}\\ \textup{universal Jacobians of type $(g,n)$}\end{Bmatrix}\hookrightarrow\begin{Bmatrix}\textup{Fine compactified universal}\\ \textup{Jacobians of type $(g,n)$}\end{Bmatrix}$
@@
Let $\mathfrak{m}^{\Gamma}:=\mathfrak{m}_{|_{\mathcal{D}(\Gamma)}}$ be the stability condition on $\Gamma$ obtained by restricting $\mathfrak{m}$ to the values $\mathfrak{m}_{(e,h,A)}$ such that $(e,h,A)\in\mathcal{D}(\Gamma)$. By Proposition 2.17, for a classical stability condition $\phi$ on $\Gamma$ to be such that $\mathfrak{m}^{\Gamma}=\mathfrak{m}^{\Gamma}(\phi)$, it needs to satisfy
$\begin{cases*}\Bigl{\lceil}\frac{2+\phi_{\{1,2,3\}}-\phi_{\{4,5,6\}}}{2}\Bigr{\rceil}=1,\\ \Bigl{\lceil}\frac{2+\phi_{\{1,5,6\}}-\phi_{\{2,3,4\}}}{2}\Bigr{\rceil}=1,\\ \Bigl{\lceil}\frac{2+\phi_{\{3,4,5\}}-\phi_{\{1,2,6\}}}{2}\Bigr{\rceil}=1,\\ \end{cases*}$
@@
$\begin{cases*}\Bigl{\lceil}\frac{2+\phi_{\{1,2,5\}}-\phi_{\{3,4,6\}}}{2}\Bigr{\rceil}=2,\\ \Bigl{\lceil}\frac{2+\phi_{\{1,3,4\}}-\phi_{\{2,5,6\}}}{2}\Bigr{\rceil}=2,\\ \Bigl{\lceil}\frac{2+\phi_{\{3,5,6\}}-\phi_{\{1,2,4\}}}{2}\Bigr{\rceil}=2.\\ \end{cases*}$
@@
$\begin{cases*}\phi_{\{1\}}-\phi_{\{2\}}+\phi_{\{3\}}-\phi_{\{4\}}+\phi_{\{5\}}-\phi_{\{6\}}<0,\\ \phi_{\{1\}}-\phi_{\{2\}}+\phi_{\{3\}}-\phi_{\{4\}}+\phi_{\{5\}}-\phi_{\{6\}}>0,\end{cases*}$
which is clearly impossible. Therefore, this example shows that there exists a fine compactified universal Jacobian such that not all of its fibres are classically polarised.
Moreover, one can extend trivially such a stability condition to obtain the same result over ¯$\overline{\mathcal{M}}_{2,n}$ for any $n\geq 6$.
@@
By work of Rhys Wells [Wel23], this example is minimal: given any curve $X$ of genus $2$ with $5$ or less marked points, every fine compactified Jacobian of $X$ is classical.
@@
One can check that it satisfies the properties of Definition 2.4. Suppose there exists a classical universal stability condition $\phi$ such that $\mathfrak{m}=\mathfrak{m}(\phi)$. It would satisfy
(3.1) $\begin{cases*}\Bigl{\lceil}\frac{2+\phi_{\{1,2\}}-\phi_{\{3,4\}}}{2}\Bigr{\rceil}=0,\\ \Bigl{\lceil}\frac{2+\phi_{\{1,3\}}-\phi_{\{2,4\}}}{2}\Bigr{\rceil}=0,\\ \lceil\phi_{\{2,3,4\}}\rceil=1,\\ \lceil\phi_{\{1,2,3\}}\rceil=3,\end{cases*}$
@@
So, we can conclude that $\mathfrak{m}$ is not induced by a classical universal stability condition. However, by remark 3.11, or simply by noticing that no single graph $\Gamma$ in $G_{2,4}$ is such that $(3,0,\{1,2\}),(3,0,\{1,3\})$, $(2,0,\{2,3,4\}),(2,0,\{1,2,3\})\in\mathcal{D}(\Gamma)$ and that no contradictions arise when removing any of the conditions in (3.1), we deduce that $\mathfrak{m}$ restricts to a classical stability condition on the dual graph of every geometric point of ¯$\overline{\mathcal{M}}_{2,4}$.
@@
A universal stability condition $\mathfrak{m}$ on $G_{2,n}$ is equivalent to a ”universal pair” $(f^{2},f^{3})$ appearing in the ”f2_f3_txt_poly_dims” repository in [Wel23], which contains computations by Rhys Wells based on a previous version of this paper. There it is shown that, for $1\leq n\leq 3$, if a universal stability condition $\mathfrak{m}$ is such that $\mathfrak{m}_{(2,0,\{1\})}\in\{0,1\}$ and $\mathfrak{m}_{(2,0,\{i\})}=0$ for each $i\neq 1$, then $\mathfrak{m}$ is induced by a classical universal stability condition. Thus, the claim follows from Lemma 3.7. ∎
Combining Examples 3.10 and 3.12, Remark 3.11 and Proposition 3.13, we can show the following result, which describes completely the inclusion of fine classical compactified Jacobians into the class of all fine compactified universal Jacobians over ¯$\overline{\mathcal{M}}_{2,n}$.
@@
$\begin{Bmatrix}\textup{Fine classical}\\ \textup{compactified universal}\\ \textup{Jacobians of type $(2,n)$}\end{Bmatrix}\hookrightarrow\begin{Bmatrix}\textup{Fine compactified universal}\\ \textup{Jacobians with classical}\\ \textup{fibres of type $(2,n)$}\end{Bmatrix}\hookrightarrow\begin{Bmatrix}\textup{Fine compactified}\\ \textup{universal Jacobians}\\ \textup{of type $(2,n)$}\end{Bmatrix}$
@@
We now treat the case of genus $g=3$ and show that there exist non-classical fine compactified universal Jacobians $\overline{\mathcal{J}}(\mathfrak{m})\subset\textup{Simp}^{d}(\overline{\mathcal{C}}_{3,n}/\overline{\mathcal{M}}_{3,n})$ if and only if $n\geq 2$.
@@
We want to show that for each universal stability condition $\mathfrak{m}$ identifying a fine compactified universal Jacobian, there exists a universal $\phi$-stability condition such that $\mathfrak{m}=\mathfrak{m}(\phi)$.
@@
It remains to show that, given any stability condition $\mathfrak{m}$ such that $\mathfrak{m}_{(2,0,\{1\})}=A$, $\mathfrak{m}_{(3,0,\emptyset)}=B_{0},$ $\mathfrak{m}_{(3,0,\{1\})}=B_{1}$ and $\mathfrak{m}_{(4,0,\emptyset)}=C$, there exists a classical stability condition $\phi$ such that $\mathfrak{m}=\mathfrak{m}(\phi)$. Since the inequalities induced by $\mathfrak{m}_{(2,1,\emptyset)}$ and $\mathfrak{m}_{(4,0,\emptyset)}$ are the same, by lemma 2.21, this amounts to showing that there exists a solution to the following system:
@@
Let $\mathfrak{m}^{\Gamma}:=\mathfrak{m}_{|_{\mathcal{D}(\Gamma)}}$ be the $c$-stability condition on $\Gamma$ obtained by restricting $\mathfrak{m}$ to the values $\mathfrak{m}_{(e,h,A)}$ such that $(e,h,A)\in\mathcal{D}(\Gamma)$. By Proposition 2.17, for a classical stability condition $\phi$ on $\Gamma$ to be such that $\mathfrak{m}^{\Gamma}=\mathfrak{m}^{\Gamma}(\phi)$, it needs to satisfy
$\begin{cases*}\Bigl{\lceil}\frac{8+2\phi_{\{1\}}-2\phi_{\{2\}}}{4}\Bigr{\rceil}-1=2;\\ \Bigl{\lceil}\frac{8+2\phi_{\{1\}}-2\phi_{\{2\}}}{4}\Bigr{\rceil}=2;\end{cases*}$
which doesn’t admit solution. Therefore $\mathfrak{m}$ defines a fine $c$-compactified universal Jacobian, whose fibres are not all classically polarised. In particular, there exists no classical universal stability condition $\phi$ such that $\mathfrak{m}=\mathfrak{m}(\phi)$.
@@
Finally, the following example shows that if the genus is at least $4$, then universal Jacobians whose fibres are not all classical always exist as long as there is at least $1$ marked point.
@@
Therefore, by Theorem A, there exists a fine $c$-compactified universal Jacobian $\overline{\mathcal{J}}(\mathfrak{m})$ over ¯$\overline{\mathcal{M}}_{g,1}$ associated to $\mathfrak{m}$.
@@
By Proposition 2.17, if $\mathfrak{m}^{\Gamma}=\mathfrak{m}^{\Gamma}(\phi)$ for any classical stability condition $\phi$ on $\Gamma$, such a $\phi$ should satisfy
$\begin{cases*}\Bigl{\lceil}\frac{4+(2g-4)\phi_{\{1\}}}{2g-2}\Bigr{\rceil}-1=\mathfrak{m}_{(4,0,\{1\})}=1;\\ \Bigl{\lceil}\frac{4+(2g-4)\phi_{\{1\}}}{2g-2}\Bigr{\rceil}=\mathfrak{m}_{(2,1,\{1\})}=1;\end{cases*}$
which has no solution. Therefore $\mathfrak{m}$ defines a fine compactified universal Jacobian, whose fibres are not all classically polarised. In particular, there exists no classical universal stability condition $\phi$ such that $\mathfrak{m}=\mathfrak{m}(\phi)$.
of Theorem 3.9.
@@
  • $g=0$.
    The statement is trivial;
  • $g=1$.
    It follows from the combination of Remark $6.13$ and Example $6.15$ of [PT22];
  • $g=2$ and $n\leq 3$.
    It is a special case of Corollary 3.14;
  • $g=3$ and $n=1$.
    It is a combination of Proposition 3.15 and Example 3.16;
  • $n=0$.
    The ”if” statement is proven in [PT23, Section 9.3], while Example 3.17 proves the other direction.
@@
Let $\Gamma$ be a graph. In [PT23, Question 8.4], the authors asked whether all stability conditions $\sigma_{\Gamma}$ on $\Gamma$ are induced by a classical stability condition $\phi$. Example 1.27 of [Viv23] shows that this is not the case.

### v2

@@ 3.4. Comparison with classical fine compactified universal Jacobians
In this section, we use our classification in terms of universal stability conditions to construct explicit examples of non (semi-)classical fine compactified universal Jacobians. Not only, we show that in general, for any $g\geq 1$ the class of fine compactified universal Jacobians is strictly larger than the class of classical fine compactified universal Jacobian described in [KP19], but we also find the minimum number of markings $n$ for this to be true over ¯$\overline{\mathcal{M}}_{g,n}$.
Moreover, we are able to make a further distinction between semi-classical fine compactified universal Jacobians and all fine compactified universal Jacobians. In particular, for genus $g=2$, with the help of computations by Rhys Wells [Wel23], we see that there is a range $4\leq n\leq 5$ where not all the fine compactified universal Jacobians are classical, but their restrictions over all geometric points are.
In the body of [PT, Section $10$] Question $\star$1 is asked. A complete answer is given in Part $1$ of Theorem B, which will be proven via the examples and propositions of this section, together with the following
@@
$\begin{Bmatrix}\textup{classical fine compactified}\\ \textup{universal Jacobians of type $(g,n)$}\end{Bmatrix}\hookrightarrow\begin{Bmatrix}\textup{Fine compactified universal}\\ \textup{Jacobians of type $(g,n)$}\end{Bmatrix}$
@@
For the whole section, we set
$\mathfrak{m}_{(1,h,A)}=\begin{cases*}0,\textup{ if }1\in A\\ d,\textup{ if }1\notin A\end{cases*},\textup{ for each }(1,h,A)\in\mathcal{D}_{g,n},$
for any universal stability condition $\mathfrak{m}$ of type $(g,n)$.
@@
Let $\mathfrak{m}^{\Gamma}:=\mathfrak{m}_{|_{\mathcal{D}(\Gamma)}}$ be the stability condition on $\Gamma$ obtained by restricting $\mathfrak{m}$ to the values $\mathfrak{m}_{(e,h,A)}$ such that $(e,h,A)\in\mathcal{D}(\Gamma)$. By Proposition 2.17, for a numerical polarisation $\phi^{\Gamma}$ on $\Gamma$ to be such that $\mathfrak{m}^{\Gamma}=\mathfrak{m}^{\Gamma}(\phi^{\Gamma})$, it needs to satisfy
$\begin{cases*}\Bigl{\lceil}\frac{2+\phi^{\Gamma}_{\{1,2,3\}}-\phi^{\Gamma}_{\{4,5,6\}}}{2}\Bigr{\rceil}=1,\\ \Bigl{\lceil}\frac{2+\phi^{\Gamma}_{\{1,5,6\}}-\phi^{\Gamma}_{\{2,3,4\}}}{2}\Bigr{\rceil}=1,\\ \Bigl{\lceil}\frac{2+\phi^{\Gamma}_{\{3,4,5\}}-\phi^{\Gamma}_{\{1,2,6\}}}{2}\Bigr{\rceil}=1,\\ \end{cases*}$
@@
$\begin{cases*}\Bigl{\lceil}\frac{2+\phi^{\Gamma}_{\{1,2,5\}}-\phi^{\Gamma}_{\{3,4,6\}}}{2}\Bigr{\rceil}=2,\\ \Bigl{\lceil}\frac{2+\phi^{\Gamma}_{\{1,3,4\}}-\phi^{\Gamma}_{\{2,5,6\}}}{2}\Bigr{\rceil}=2,\\ \Bigl{\lceil}\frac{2+\phi^{\Gamma}_{\{3,5,6\}}-\phi^{\Gamma}_{\{1,2,4\}}}{2}\Bigr{\rceil}=2.\\ \end{cases*}$
@@
$\begin{cases*}\phi^{\Gamma}_{\{1\}}-\phi^{\Gamma}_{\{2\}}+\phi^{\Gamma}_{\{3\}}-\phi^{\Gamma}_{\{4\}}+\phi^{\Gamma}_{\{5\}}-\phi^{\Gamma}_{\{6\}}<0,\\ \phi^{\Gamma}_{\{1\}}-\phi^{\Gamma}_{\{2\}}+\phi^{\Gamma}_{\{3\}}-\phi^{\Gamma}_{\{4\}}+\phi^{\Gamma}_{\{5\}}-\phi^{\Gamma}_{\{6\}}>0,\end{cases*}$
which is clearly impossible. Therefore, this example shows that there exists a non semi-classical fine compactified universal Jacobian.
Moreover, one can extend the stability condition $\mathfrak{m}$ to $\mathfrak{m}^{\prime}$ over ¯$\overline{\mathcal{M}}_{2,n}$ for any $n\geq 6$, by imposing $\mathfrak{m}_{(e,h,A)}^{\prime}:=\mathfrak{m}_{(e,h,A\cap[6])}$. Thus, the same result holds for any $n\geq 6$.
@@
By work of Rhys Wells [Wel23], this example is minimal: given any $n$-pointed curve $X$ of genus $2$ with $n\leq 5$, every fine compactified Jacobians of $X$ is classical.
@@
One can check that it satisfies the properties of Definition 2.4. Suppose there exists a universal numerical polarisation $\phi$ such that $\mathfrak{m}=\mathfrak{m}(\phi)$. It would satisfy
(3.1) $\begin{cases*}\Bigl{\lceil}\frac{-1+\phi_{\{1,2\}}-\phi_{\{3,4\}}}{2}\Bigr{\rceil}=0,\\ \Bigl{\lceil}\frac{-1+\phi_{\{1,3\}}-\phi_{\{2,4\}}}{2}\Bigr{\rceil}=0,\\ \lceil\phi_{\{2,3,4\}}\rceil=1,\\ \lceil\phi_{\{1,2,3\}}\rceil=3,\end{cases*}$
@@
So, we can conclude that $\mathfrak{m}$ is not a classical universal stability condition. However, by remark 3.11, or simply by noticing that no single graph $\Gamma$ in $G_{2,4}$ is such that $(3,0,\{1,2\}),(3,0,\{1,3\})$, $(2,0,\{2,3,4\}),(2,0,\{1,2,3\})\in\mathcal{D}(\Gamma)$ and that no contradictions arise when removing any of the conditions in (3.1), we deduce that $\mathfrak{m}$ restricts to a classical stability condition on the dual graph of every geometric point of ¯$\overline{\mathcal{M}}_{2,4}$, hence, it is semi-classical.
@@
A universal stability condition $\mathfrak{m}$ on $G_{2,n}$ is equivalent to a ”universal pair” $(f^{2},f^{3})$ appearing in the ”f2_f3_txt_poly_dims” repository in [Wel23], which contains computations by Rhys Wells based on a previous version of this paper. There, it is shown that, for $1\leq n\leq 3$, if a universal stability condition $\mathfrak{m}$ is such that $\mathfrak{m}_{(2,0,\{1\})}\in\{0,1\}$ and $\mathfrak{m}_{(2,0,\{i\})}=0$ for each $i\neq 1$, then $\mathfrak{m}$ is classical. Thus, the claim follows from Lemma 3.7. ∎
Combining Examples 3.10 and 3.12, Remark 3.11 and Proposition 3.13, we can show the following result, which describes completely the inclusion of fine (semi-)classical compactified Jacobians into the class of all fine compactified universal Jacobians over ¯$\overline{\mathcal{M}}_{2,n}$.
@@
$\begin{Bmatrix}\textup{classical fine}\\ \textup{compactified universal}\\ \textup{Jacobians of type $(2,n)$}\end{Bmatrix}\hookrightarrow\begin{Bmatrix}\textup{semi-classical fine}\\ \textup{compactified universal}\\ \textup{Jacobians of type $(2,n)$}\end{Bmatrix}\hookrightarrow\begin{Bmatrix}\textup{Fine compactified}\\ \textup{universal Jacobians}\\ \textup{of type $(2,n)$}\end{Bmatrix}$
@@
We now treat the case of genus $g=3$ and show that there exist non semi-classical fine compactified universal Jacobians $\overline{\mathcal{J}}(\mathfrak{m})\subset\textup{Simp}^{d}(\overline{\mathcal{C}}_{3,n}/\overline{\mathcal{M}}_{3,n})$ if and only if $n\geq 2$.
@@
We want to show that for each universal stability condition $\mathfrak{m}$ identifying a fine compactified universal Jacobian, there exists a universal numerical polarisation $\phi$ such that $\mathfrak{m}=\mathfrak{m}(\phi)$.
@@
It remains to show that, given any stability condition $\mathfrak{m}$ such that $\mathfrak{m}_{(2,0,\{1\})}=A$, $\mathfrak{m}_{(3,0,\emptyset)}=B_{0},$ $\mathfrak{m}_{(3,0,\{1\})}=B_{1}$ and $\mathfrak{m}_{(4,0,\emptyset)}=C$, there exists a universal numerical polarisation $\phi$ such that $\mathfrak{m}=\mathfrak{m}(\phi)$. Since the inequalities induced by $\mathfrak{m}_{(2,1,\emptyset)}$ and $\mathfrak{m}_{(4,0,\emptyset)}$ are the same, by lemma 2.21, this amounts to showing that there exists a solution to the following system:
@@
Let $\mathfrak{m}^{\Gamma}:=\mathfrak{m}_{|_{\mathcal{D}(\Gamma)}}$ be the $c$-stability condition on $\Gamma$ obtained by restricting $\mathfrak{m}$ to the values $\mathfrak{m}_{(e,h,A)}$ such that $(e,h,A)\in\mathcal{D}(\Gamma)$. By Proposition 2.17, for a numerical polarisation $\phi^{\Gamma}$ on $\Gamma$ to be such that $\mathfrak{m}^{\Gamma}=\mathfrak{m}^{\Gamma}(\phi^{\Gamma})$, it needs to satisfy
$\begin{cases*}\Bigl{\lceil}\frac{8+2\phi^{\Gamma}_{\{1\}}-2\phi^{\Gamma}_{\{2\}}}{4}\Bigr{\rceil}-1=2;\\ \Bigl{\lceil}\frac{8+2\phi^{\Gamma}_{\{1\}}-2\phi^{\Gamma}_{\{2\}}}{4}\Bigr{\rceil}=2;\end{cases*}$
which doesn’t admit solution. Therefore $\mathfrak{m}$ defines a fine non semi-classical compactified universal Jacobian. In particular, there exists no universal numerical polarisation $\phi$ such that $\mathfrak{m}=\mathfrak{m}(\phi)$.
@@
Finally, the following example shows that if the genus is at least $4$, then universal Jacobians that are non semi-classical always exist as long as there is at least $1$ marked point.
@@
Therefore, by Theorem A, there exists a fine compactified universal Jacobian $\overline{\mathcal{J}}(\mathfrak{m})$ over ¯$\overline{\mathcal{M}}_{g,1}$ associated to $\mathfrak{m}$.
@@
By Proposition 2.17, if $\mathfrak{m}^{\Gamma}=\mathfrak{m}^{\Gamma}(\phi^{\Gamma})$ for any numerical polarisation $\phi^{\Gamma}$ on $\Gamma$, such a $\phi^{\Gamma}$ should satisfy
$\begin{cases*}\Bigl{\lceil}\frac{4+(2g-4)\phi^{\Gamma}_{\{1\}}}{2g-2}\Bigr{\rceil}-1=\mathfrak{m}_{(4,0,\{1\})}=1;\\ \Bigl{\lceil}\frac{4+(2g-4)\phi^{\Gamma}_{\{1\}}}{2g-2}\Bigr{\rceil}=\mathfrak{m}_{(2,1,\{1\})}=1;\end{cases*}$
which has no solution. Therefore $\mathfrak{m}$ defines a fine non semi-classical compactified universal Jacobian. In particular, there exists no universal numerical polarisation $\phi$ such that $\mathfrak{m}=\mathfrak{m}(\phi)$.
Proof of Theorem 3.9.
@@
  • $g=0$. The statement is trivial;
  • $g=1$. It follows from the combination of Remark $6.13$ and Example $6.15$ of [PT23];
  • $g=2$ and $n\leq 3$. It is a special case of Corollary 3.14;
  • $g=3$ and $n=1$. It is a combination of Proposition 3.15 and Example 3.16;
  • $n=0$. The ”if” statement is proven in [PT, Section 9.3], while Example 3.17 proves the other direction.
@@
Let $\Gamma$ be a graph. In [PT, Question 8.4], the authors asked whether all stability conditions $\sigma_{\Gamma}$ on $\Gamma$ are induced by a numerical polarisation $\phi^{\Gamma}$. Example 1.27 of [Viv] shows that this is not the case.